OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−3, 2) and (2, 1). y + 3 = −5(x − 2) y − 2 = −5(x + 3) y + 3 = −(x − 2) y − 2 = −(x + 3)

OpenStudy (blurbendy):

point slope form = y - y1 = slope(x - x1)

OpenStudy (anonymous):

ok

OpenStudy (blurbendy):

i would recommend finding the slope first, given the two points you have

OpenStudy (anonymous):

which is - 1/5

OpenStudy (anonymous):

so its c but C is -1/5

OpenStudy (blurbendy):

? your c just says -

OpenStudy (blurbendy):

not -1/5

OpenStudy (anonymous):

i know but that was a typo

OpenStudy (blurbendy):

what does d say then?

OpenStudy (anonymous):

C and D are both -1/5

OpenStudy (anonymous):

the difference is D is +

OpenStudy (blurbendy):

okay, so now choose the first point, and substitute into the formula i gave you

OpenStudy (anonymous):

which turns into C :)

OpenStudy (blurbendy):

no, it turns into D y1 = 2 x1 = -3 y - 2 = -1/5(x -(-3))

OpenStudy (blurbendy):

understand?

OpenStudy (anonymous):

yes :) so c is correct. thank u :) mind helping with 2 more?

OpenStudy (anonymous):

i mean D

OpenStudy (blurbendy):

no, d is correct

OpenStudy (blurbendy):

good :)

OpenStudy (anonymous):

opps sorry meant D :)

OpenStudy (anonymous):

so mind with 2 more?

OpenStudy (blurbendy):

ok

OpenStudy (anonymous):

Given the equation y − 3 = (x + 6) in point-slope form, identify the equation of the same line in slope-intercept form. x − 2y = −12 y = 1/2x y = 1/2x + 6 y = 1/2x + 9

OpenStudy (blurbendy):

slope intercept form = y = mx + b where m is the slope

OpenStudy (blurbendy):

did you forget to write the slope?

OpenStudy (anonymous):

opps yes hold on

OpenStudy (anonymous):

slope is 1/2

OpenStudy (blurbendy):

okay, so put y - 3 = 1/2(x + 6) into y = mx + b form (hint, first distribute the 1/2)

OpenStudy (anonymous):

so i got y-3 = 1/2x + 3

OpenStudy (blurbendy):

yup, keep going

OpenStudy (anonymous):

so the answer would be C

OpenStudy (blurbendy):

oh, nevermind

OpenStudy (anonymous):

? is that correct?

OpenStudy (blurbendy):

you got it

OpenStudy (anonymous):

thxs! :) mind one more?

OpenStudy (blurbendy):

sure

OpenStudy (anonymous):

Given the equation y − 4 =3/4 (x + 8) in point-slope form, identify the equation of the same line in standard form.

OpenStudy (blurbendy):

standard form = Ax + By = C

OpenStudy (blurbendy):

start by distributing the 1/2 again

OpenStudy (anonymous):

−3/4x + y = 10 3x − 4y = −40 y = 3/4x + 12 y = 3/4x + 10

OpenStudy (blurbendy):

i mean 3/4

OpenStudy (blurbendy):

not 1/2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so i got y-4 = 3/4x + 6

OpenStudy (blurbendy):

yes, and so we need it in the form: Ax + By = C

OpenStudy (anonymous):

y = 3/4x + 10?

OpenStudy (blurbendy):

no, we need x and y on the same side

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

-3/4x + y = 10?

OpenStudy (blurbendy):

perfect, just subtract the -3/4x

OpenStudy (anonymous):

so that would be the answer?

OpenStudy (blurbendy):

yes

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

aww i got it wrong :(

OpenStudy (blurbendy):

-3/4x + y = 10 was wrong? i guess they wanted you to multiply the 4 out.... lame. the question should have specified that

OpenStudy (blurbendy):

sorry :(

OpenStudy (blurbendy):

well i guess A shouldnt have been negative